# Extracting self-similarity from data

### v2

@@ Results
where $u(x,t)$ is the fluid velocity and $x$ and $t$ the spatial and temporal coordinates, respectively. The subscript denotes partial differentiation. Using the method of characteristics, the solution can be expressed as $u(x,t)=f(\xi)=f(x-ut)=f(x+\alpha)$, i.e., the problem described by Eq. (5) is invariant under a non-uniform translational transformation. In this work, we consider a sinusoidal initial condition, $u(x,0)=\sin{(2\pi x)}$, with $x\in\left[0,1\right]$. Figures 2(a) and (b) show the spatiotemporal evolution of velocity and profiles extracted at five different time instants, respectively, illustrating the steepening of the solution as time evolves. Figure 2(c) shows the algorithmically collapsed self-similar velocity profiles on the transformed coordinates $\left(\tilde{x},t\right)$, following a transformation of the form $\tilde{x}=x+\alpha(x,t)$ (simpler transformations do not succeed at collapsing the data). Since this is an initial value problem, we only consider the distance of the profiles with respect to the initial condition, i.e., the objective function in Eq. (Step 1. Search for similarity variables.) is simplified to $\sum_{i=1,\dots,n_{t}}\|u(\tilde{x},t_{i})-u(x,0)\|_{2}^{2}$. We use PySR to interpret the algorithmically computed transformation $\alpha$. The library of candidate variables is composed of the problem’s independent and dependent variables, i.e., $\alpha=\alpha(u,x,t)$. The method interprets the identified transformation as $\alpha=-ut$, which is identical to the analytical similarity transformation.
@@
We consider the statistically stationary flow past a slender bluff body at high Reynolds numbers, which is a characteristic example of a free (i.e., unconfined) turbulent flow. The term slender signifies an effectively infinite body length in one crossflow direction, along which the flow statistics can be considered homogeneous. It can be shown  (?) that such flows can be modelled using a boundary layer approximation, similarly to the one employed by Prandtl for laminar boundary layers. We note, however, that this approximation is now applied to the time-averaged flow, and that Reynolds stresses (i.e., turbulent momentum transport) are significant, i.e.,
@@
where the overbar denotes time-averaging and $\nu$ is the kinematic viscosity of the fluid. $u$ and $v$ are the flow velocities along the streamwise $x$ and crosswise $y$ directions, which are decomposed into time-averaged $(\bar{u},\bar{v})$ and fluctuating $(u^{\prime},v^{\prime})$ components. The formulation leading to Eq. (6) introduces the Reynolds stress term $\overline{u^{\prime}v^{\prime}}$ into the flow governing equations, i.e., an additional unknown which cannot be calculated implicitly (this is the well-known closure problem of turbulence). The identification of similarity in this case is, therefore, not possible from the equations themselves, unless an assumption is made regarding the relation of the unknown Reynolds stress terms to the mean velocity distribution (i.e., turbulence modelling). However, it is customary to hypothesize that the turbulent boundary layer equations accept self-similar solutions far from initial conditions, a fact also supported by experimental observations  (?, ?). This approximate self-similarity is the origin behind various scaling laws for the evolution of free shear flows, widely used in many applications of the energy and transportation sectors, such as wind farm planning  (?) and jet engine noise prediction  (?). Returning to the particular example of the slender bluff body wake, we may consider the time-averaged velocity deficit $\zeta(x,y)=U_{\infty}-\bar{u}(x,y)$ that the body produces far (i.e., tens of characteristic body lengths) downstream. In the above, $U_{\infty}$ is the constant free stream velocity. The flow statistics are homogeneous along the spanwise direction $z$. Past works  (?, ?) indicate a self-similarity of the form
@@
where $\bar{u}_{\text{cntr}}=\bar{u}(x,0)$ denotes the centreline velocity and $y_{1/2}(x)$ is the wake half-width defined such that $\bar{u}(x,\pm y_{1/2})=0.5\left(U_{\infty}+u_{\text{cntr}}\right)$. We attempt to extract the self-similar relation Eq. (7) directly from experimental data. To this end, we measured the turbulent wake of a plate of 53% porosity immersed in a water flume normal to the flow (see Fig. 3(a)) at a Reynolds number based on the free stream velocity and plate width $Re\approx 6,000$. The velocity fields at various positions downstream of the plate were measured using Particle Image Velocimetry (Phantom 4MP camera at 50 Hz acquisition frequency) as shown in Fig. 3(a). Each mean streamwise velocity profile was then calculated by averaging 3,000 vector fields. More information regarding the experimental procedure can be found in  (?). We consider the mean velocities at five stations in the wake of the plate (Fig. 3(b)). Figure 3(c) shows the collapse that is obtained via the proposed method, assuming transformations of the form $\tilde{y}=\alpha(x)y$ and $\tilde{u}=\beta(x)\bar{u}+\gamma(x)$ (simpler transformations cannot collapse the data). The identified transformations are regressed as functions of the characteristic scales and variables of the problem using PySR, with $\alpha=\alpha(x,y_{w},y_{1/2})$ (with $y_{w}(x)$ defined as $\bar{u}(x,\pm y_{w})=0.99U_{\infty}$), $\beta=\beta(U_{\infty},\bar{u}_{\text{cntr}},x,\nu)$ and $\gamma=\gamma(U_{\infty},\bar{u}_{\text{cntr}},x,\nu).$ The interpreted similarity transformations are
@@
which are close (within an arbitrary multiplicative constant) to the self-similar expressions for turbulent wakes found in the literature  (?, ?, ?). Besides our own experiments, we also test our method by using experimental data available in the literature  (?), and again retrieve the self-similarity relations for turbulent wakes (see Supplementary Material).
@@
The bursting of bubbles at the sea surface plays a crucial role in the exchanges between oceans and the atmosphere, thereby being important for climate and weather  (?). When a bubble bursts at the sea surface, aerosol is produced via two mechanisms: the rupture of the bubble’s cap film, and the formation of a vertical jet that breaks into droplets following cavity collapse  (?). The jet and aerosol properties are a consequence of the non-linear fluid dynamics near the points where topology changes  (?, ?). In the case of a collapsing cavity, several studies have found that near the collapse of the travelling capillary waves at the axis of symmetry, the liquid-gas interface evolves in a self-similar manner, enabling the derivation of scaling laws for the bubble and jet dynamics  (?, ?, ?, ?, ?, ?).
@@
We simulate the collapse of a cavity ($\text{Bo}=10^{-3}$, ${\text{La}}=2500$) by solving the two-phase incompressible axisymmetric Navier-Stokes equations using Basilisk (www.basilisk.fr), which is an open-source flow solver that has been used in computational studies of bursting bubbles  (?, ?, ?). The evolution of the liquid-gas interface is shown in Figs. 4(a-d). We extract profiles of the interface $h(r,t)$ as it approaches the axis of symmetry, at $t_{*}=\left(t_{0}-t\right)/t_{c}=\left[10,9,8,7,6\right]$, where $t_{c}$ is the characteristic time of the horizontal capillary wave  (?), and $t_{0}$ is the moment where the capillary waves meet at the axis of symmetry, $r=0$. The extracted profiles are shown in Fig. 4(e).  (?, ?) observed that in this time window, the interface approximately follows the $\left(t_{0}-t\right)^{2/3}$ scaling of inviscid theory  (?). The algorithm successfully collapses the extracted profiles into a single curve (Fig. 4(f)) for transformations of the form $\tilde{r}=\alpha(t)r$ for the radial coordinate and $\tilde{h}=\beta(t)h+\gamma(t)$ for the liquid-gas interface. By regressing the identified transformations $\alpha$, $\beta$, and $\gamma$ as power-law functions of the non-dimensional time $t_{*}$ (Fig. 4(g-i)), we retrieve scalings close (apart from an arbitrary multiplicative constant) to the theoretically derived ones  (?, ?, ?, ?, ?).
Grid turbulence
@@
We consider the case of homogeneous decaying turbulence, experimentally realized by passing a stream of fluid through a uniformly spaced grid inside a wind tunnel (Fig. 5(a)). We analyze measurements of velocity time-series, obtained via hot-wire anemometry at seven positions downstream of the grid (the details of the experimental campaign can be found in Ref.  (?)). The measured turbulence is fully developed, approximately homogeneous and yields the -5/3 law for the energy spectrum at intermediately-sized eddies, as predicted by Kolmogorov’s K41 framework  (?, ?, ?). This configuration puts our algorithm to a real test: Our algorithm can identify self-similarities based on a single scale, while the turbulence cascade is generally accepted to be a two-scale problem.
@@
The theoretical derivation of the -5/3 law assumes that at sufficiently high Reynolds numbers, an intermediate self-similar region forms in the cascade, which is independent of large- and small scale effects  (?, ?, ?). However, the above description is part of a more complex flow-picture, as the widely studied paradigm of grid turbulence sufficiently far from initial conditions shows. The turbulence cascade in that case is generally accepted to be characterized by two self-similarities, both present at the same time and at different eddy sizes: One at large scales where viscosity is negligible  (?, ?), and one at small scales where non-equilibrium effects are negligible  (?, ?). The energy spectrum thus accepts the following general expression  (?)
@@
where $x$, $\kappa$, $E_{11}(\kappa,x)$, $\epsilon(x)$ represent the distance from the grid, wavenumber, energy spectrum and dissipation rate, respectively. $L$ and $\eta$ are the integral and Kolmogorov scales, characteristic of the large- and small-scale self-similarities, respectively. To retrieve the -5/3 law, $E_{11}(k,x)=\epsilon(x)^{2/3}\kappa^{-5/3}$, one needs to be asymptotically far from large scales (i.e., $\kappa L\to\infty$) and far from small scales ($\kappa\eta\to 0$) at the same time, as in that case K41 assures that both $f$ and $g$ tend to unity. It is of interest to see how our single-scale algorithm fares in this two-scale problem.
Figure 5(b) plots the measured power spectral densities versus the wavenumbers at different measurement stations. Figures 5(c,d) show the collapse of the power spectral densities using large and small scale similarity variables, respectively. Using a library that consists of $(L,\eta)$ for the $x$ axis and $(\epsilon,L,\eta,k)$ for the $y$ axis, symbolic regression of the algorithmically identified stretching transformations yields the following expressions, plotted in Fig. 5(e)
@@
As expected, our algorithm fails to derive the two-scale similarity expression (9). Further insight may be obtained by using the dissipation scaling $\epsilon\propto k^{3/2}/L$, where $k$ is the turbulence kinetic energy, which is known to characterize homogeneous decaying turbulence far from initial conditions  (?, ?, ?). Expression (10) then becomes
@@
Inspection of the above expression reveals that our algorithm has produced novel information by extracting self-similarity based on a single empirical length scale, $l\approx L^{0.38}\eta^{0.62}$, which is a combination of $L$ and $\eta$. It is very interesting that a single length scale theory for the cascade does in fact exist (see Ref.  (?)). The single length scale of that theory is the so-called Taylor microscale, which in our problem becomes $\lambda=L^{1/3}\eta^{2/3}$  (?), i.e., very close to our empirically-derived length scale. Therefore, while our algorithm may not be appropriate for the extraction of two-scale self-similarities, it is in fact able to contribute to the discussion regarding the modeling and behaviour of the turbulence cascade.
@@
Figure 5: Data-driven identification of self-similarity in decaying turbulence. (a) Schematic of the experimental set-up. (b) Experimentally measured power spectral densities. The dashed vertical line delineates the range of the spectrum that is used. (c) Measured spectrum normalised by the inertial scales. (d) Measured spectrum normalised by the Kolmogorov scales. (e) Measured spectrum normalised by the algorithmically identified expression (Eq. (10)).
Discussion
We have presented a methodology for identifying similarity variables from data, in the absence of governing equations. Given measurements of a quantity of interest, the method computes the discrete values of the similarity transformations that best collapse the data and expresses their analytic form via symbolic regression. The self-similarity that we identify is only approximate, in the sense that it is the best one that fits the data, and is not derived from the governing partial differential equations and the boundary conditions of the specific problem. The capabilities of the method were demonstrated in five fluid mechanical problems, which are known or assumed to accept similarity solutions, based on data from both numerical and laboratory experiments. In four problems, the transformations identified by the method were in agreement to the analytically or empirically derived similarity transformations, while circumventing the need for scaling analysis or similarity arguments. In the fifth problem (turbulence cascade) our method derived a novel expression for the energy spectrum, different from the widely accepted two-scale similarity expressions, contributing in that way to the discussion regarding the nature of similarity in turbulence and showcasing how the present work opens up opportunities for identification of similarity and scaling laws in situations where rigorous mathematical analysis is challenging. This includes a wide range of applications in fluid physics but also in other processes such as quasicrystal shape and growth  (?), stellar collapse  (?), single protein dynamics  (?), and others.
Acknowledgments

### v3

@@ 3 Results
where $u(x,t)$ is the fluid velocity and $x$ and $t$ the spatial and temporal coordinates, respectively. The subscript denotes partial differentiation. Using the method of characteristics, the solution can be expressed as $u(x,t)=f(\xi)=f(x-ut)=f(x+\alpha)$, i.e., the problem described by Eq. (7) is invariant under a non-uniform translational transformation. In this work, we consider a sinusoidal initial condition, $u(x,0)=\sin{(2\pi x)}$, with $x\in\left[0,1\right]$. Figures 2(a) and (b) show the spatiotemporal evolution of velocity and profiles extracted at seven different time instants, respectively, illustrating the steepening of the solution as time evolves. The times shown are non-dimensionalised with the time of shock formation $_{c}, t=t$$^{*}=t/t_{c}=2\pi t$. The algorithm is provided with the data shown with markers in figure 2(b).
Figure 2(c) shows the algorithmically collapsed velocity profiles on the transformed coordinates $\left(\tilde{x},t\right)$, i.e. the output of Step 1 of the algorithm, following a transformation of the form $\tilde{x}=x+\alpha(x,t)$ (simpler transformations do not succeed at collapsing the data). Since this is an initial value problem, we only consider the distance of the profiles with respect to the initial condition, i.e., the objective function in Eq. (2) is simplified to $\sum_{i=1,\dots,n_{t}}\|u(\tilde{x},t_{i})-u(x,0)\|_{2}^{2}$.
We use PySR to interpret the algorithmically computed transformation $\alpha$. The library of candidate variables is composed of the problem’s independent and dependent variables, i.e., $\alpha=\alpha(u,x,t)$, and the library of operators includes the four basic mathematical operations ($+,-,\times,\div$). PySR interprets the algorithmically computed transformation as $\alpha=-ut$, which is identical to the analytical similarity transformation. Close inspection of figure 2(c) shows that the profiles are not perfectly collapsed (for example, at $t^{*}=0.936$). The range-normalised mean absolute error between the algorithmically computed discrete values of the transformation and the analytical ones is $2.2\%$. However, in Step 2 of the workflow (symbolic regression), the algorithm attempts to fit expressions that balance complexity and accuracy, allowing it to recover the exact analytical transformation $\alpha=-ut$.
@@
As a third example, we consider the statistically stationary flow past a slender bluff body at high Reynolds numbers, which is a characteristic example of a free (i.e., unconfined) turbulent shear flow. The term slender signifies an effectively infinite body length in one crossflow direction, along which the flow statistics can be considered homogeneous. Such flows can be modelled using a boundary layer approximation, similar to the one employed by Prandtl for laminar boundary layers, applied to the time-averaged flow (Townsend, 1976), i.e.,
@@
where the overbar denotes time-averaging and $\nu$ is the kinematic viscosity of the fluid. $u$ and $v$ are the flow velocities along the streamwise $x$ and crosswise $y$ directions, which are decomposed into time-averaged $(\bar{u},\bar{v})$ and fluctuating $(u^{\prime},v^{\prime})$ components. The formulation leading to Eq. (8) introduces the Reynolds stress term $\overline{u^{\prime}v^{\prime}}$ into the flow governing equations, i.e., an additional unknown which cannot be calculated implicitly (this is the well-known closure problem of turbulence). The identification of similarity in this case is, therefore, not possible from the equations themselves, unless an assumption is made regarding the relation of the unknown Reynolds stress terms to the mean velocity distribution (i.e., turbulence modelling). However, it is customary to hypothesize that the turbulent boundary layer equations accept self-similar solutions far from initial conditions, a fact also supported by experimental observations (Townsend, 1976; Cantwell, 2002). This approximate self-similarity is the origin behind various scaling laws for the evolution of free shear flows, widely used in many applications of the energy and transportation sectors, such as wind farm planning (Bempedelis & Steiros, 2022) and jet engine noise prediction (Tam, 2019). Returning to the particular example of the slender bluff body wake, we may consider the time-averaged velocity deficit $\zeta(x,y)=U_{\infty}-\bar{u}(x,y)$ that the body produces far (i.e., tens of characteristic body lengths) downstream. In the above, $U_{\infty}$ is the constant free stream velocity. The flow statistics are homogeneous along the spanwise direction $z$. Wake self-similarity is then known to assume the form (Pope, 2000; Townsend, 1976)
@@
where $\bar{u}_{\text{cntr}}=\bar{u}(x,0)$ denotes the centreline velocity and $y_{1/2}(x)$ is the wake half-width defined such that $\bar{u}(x,\pm y_{1/2})=0.5\left(U_{\infty}+u_{\text{cntr}}\right)$. We attempt to extract the self-similar relation Eq. (9) directly from experimental data. To this end, we measured the turbulent wake of a plate of 53% porosity immersed in a water flume normal to the flow (see figure 8(a)) at a Reynolds number based on the free stream velocity and plate width $Re\approx 6,000$. The velocity fields at various positions downstream of the plate were measured using Particle Image Velocimetry (Phantom 4MP camera at 50 Hz acquisition frequency) as shown in figure 8(a). Each mean streamwise velocity profile was then calculated by averaging 3,000 vector fields. More information regarding the experimental procedure can be found in Bekoglu et al. (2024, 2025). We consider the mean velocities at five stations in the wake of the plate (figure 8(b)). Figure 8(c) shows the collapse that is obtained via the proposed method, assuming transformations of the form $\tilde{y}=\alpha(x)y$ and $\tilde{u}=\beta(x)\bar{u}+\gamma(x)$ (simpler transformations cannot collapse the data). The identified transformations are regressed as functions of the characteristic scales and variables of the problem using PySR, with $\alpha=\alpha(x,y_{w},y_{1/2})$ (with $y_{w}(x)$ defined as $\bar{u}(x,\pm y_{w})=0.99U_{\infty}$), $\beta=\beta(U_{\infty},\bar{u}_{\text{cntr}},x,\nu)$ and $\gamma=\gamma(U_{\infty},\bar{u}_{\text{cntr}},x,\nu)$. The library of operators consists of the four basic mathematical operations ($+,-,\times,\div$). The interpreted similarity transformations are
@@
which match (ignoring the arbitrary multiplicative constant) the self-similar expressions for turbulent wakes found in the literature (Tennekes & Lumley, 1972; Pope, 2000; Townsend, 1976). Besides our own experiments, we also test our method by using experimental data available in the literature (Cimbala et al., 1988), and again retrieve the self-similarity relations for turbulent wakes (see Appendix C).
@@
The bursting of bubbles at the sea surface plays a crucial role in the exchanges between oceans and the atmosphere, thereby being important for climate and weather (Deike, 2022). When a bubble bursts at the sea surface, aerosol is produced via two mechanisms: the rupture of the bubble’s cap film, and the formation of a vertical jet that breaks into droplets following cavity collapse (Deike, 2022). The jet and aerosol properties are a consequence of the non-linear fluid dynamics near the points where topology changes (Eggers, 1997; Eggers & Fontelos, 2015). In the case of a collapsing cavity, several studies have found that near the collapse of the travelling capillary waves at the axis of symmetry, the liquid-gas interface evolves in a self-similar manner, enabling the derivation of scaling laws for the bubble and jet dynamics (Keller & Miksis, 1983; Zeff et al., 2000; Duchemin et al., 2002; Ghabache et al., 2014; Gañán-Calvo, 2017; Deike et al., 2018).
@@
We simulate the collapse of a cavity ($\text{Bo}=\rho gR_{0}^{2}/\gamma=10^{-3}$, ${\text{La}}=\rho\gamma R_{0}/\mu^{2}=2500$) by solving the two-phase incompressible axisymmetric Navier-Stokes equations using Basilisk (www.basilisk.fr), an open-source flow solver that has been previously used in computational studies of bursting bubbles (Lai et al., 2018; Berny et al., 2020; Sanjay et al., 2021). In the above, $\rho$ and $\mu$ are the liquid density and viscosity, respectively, $\gamma$ is the interfacial tension, $R_{0}$ is the initial bubble radius, and $g$ is the gravitational acceleration. The evolution of the liquid-gas interface is shown in figures 4(a-d). We extract profiles of the interface $h(r,t)$ as it approaches the axis of symmetry, at $t_{*}=\left(t_{0}-t\right)/t_{c}=\left[10,9,8,7,6\right]$, where $t_{c}$ is the characteristic time of the horizontal capillary wave (Lai et al., 2018), and $t_{0}$ is the moment where the capillary waves meet at the axis of symmetry, $r=0$. The extracted profiles are shown in figure 4(e). Duchemin et al. (2002); Lai et al. (2018) observed that in this time window, the interface approximately follows the $\left(t_{0}-t\right)^{2/3}$ scaling of inviscid theory (Keller & Miksis, 1983). The algorithm successfully collapses the extracted profiles into a single curve (figure 4(f)) for transformations of the form $\tilde{r}=\alpha(t)r$ for the radial coordinate and $\tilde{h}=\beta(t)h+\gamma(t)$ for the liquid-gas interface. By regressing the identified transformations $\alpha$, $\beta$, and $\gamma$ as power-law functions of the non-dimensional time $t_{*}$ (figure 4(g-i)), we retrieve the expressions
$\tilde{r}=c_{1}rt_{*}^{-0.61}$ (11a)
$\tilde{h}=c_{2}ht_{*}^{-0.69}-c_{3}h_{0}t_{*}^{-0.66}$ (11b)
which are close (ignoring the arbitrary multiplicative constants $c_{1}$, $c_{2}$, and $c_{3}$) to the theoretically derived ones (Keller & Miksis, 1983; Zeff et al., 2000; Duchemin et al., 2002; Ghabache et al., 2014; Lai et al., 2018).
Decaying turbulence
Figure 5: Data-driven identification of self-similarity in decaying turbulence. (a) Schematic of the experimental set-up. (b) Experimentally measured power spectral densities. The dashed vertical line delineates the range of the spectrum that is used. (c) Measured spectrum normalized by the inertial scales. (d) Measured spectrum normalized by the Kolmogorov scales. (e) Measured spectrum normalized by the algorithmically identified expression (Eq. (13)).
@@
We consider the case of homogeneous decaying turbulence, experimentally realized by passing a stream of fluid through a uniformly spaced grid inside a wind tunnel (figure 5(a)). We analyze measurements of velocity time-series, obtained via hot-wire anemometry at seven positions downstream of the grid (the details of the experimental campaign can be found in Steiros (2022a)). The measured turbulence is fully developed, approximately homogeneous and yields the -5/3 law for the energy spectrum at intermediately-sized eddies, as predicted by Kolmogorov’s K41 framework (Kolmogorov, 1941a, c, b).
@@
The theoretical derivation of the -5/3 law assumes that at sufficiently high Reynolds numbers, an intermediate self-similar region forms in the cascade, which is independent of large- and small scale effects (Pope, 2000; Batchelor, 1953; Frisch, 1995). However, the above description is part of a more complex flow-picture, as the widely studied paradigm of grid turbulence sufficiently far from initial conditions shows. The turbulence cascade in that case is generally accepted to be characterized by two self-similarities, both present at the same time and at different eddy sizes: One at large scales where viscosity is negligible (Steiros, 2022a; Lundgren, 2003), and one at small scales where non-equilibrium effects are negligible (Lundgren, 2003; Pope, 2000). The energy spectrum thus accepts the following general expression (Pope, 2000)
@@
where $x$, $\kappa$, $E_{11}(\kappa,x)$, $\epsilon(x)$ represent the distance from the grid, wavenumber, energy spectrum and dissipation rate, respectively. $L$ and $\eta$ are the integral and Kolmogorov scales, characteristic of the large- and small-scale self-similarities, respectively. To retrieve the -5/3 law, $E_{11}(k,x)=\epsilon(x)^{2/3}\kappa^{-5/3}$, one needs to be asymptotically far from large scales (i.e., $\kappa L\to\infty$) and far from small scales ($\kappa\eta\to 0$) at the same time, as in that case K41 assures that both $f$ and $g$ tend to unity. It is of interest to see how our algorithm fares in this two-scale problem.
Figure 5(b) plots the measured power spectral densities versus the wavenumbers at different measurement stations. Figures 5(c,d) show the collapse of the power spectral densities using large and small scale similarity variables, respectively. Using a library that consists of $(L,\eta)$ for the $x$ axis and $(\epsilon,L,\eta,k)$ for the $y$ axis, symbolic regression of the algorithmically identified dilational transformations ($\tilde{\kappa}=\alpha(x)\kappa$, $\tilde{E_{11}}=\beta(x)E_{11}$) yields the following expressions, plotted in figure 5(e)
@@
As expected, the algorithm cannot derive the two-scale similarity expression (12), as the assumed dilational transformations effectively reduce it to a single-scale algorithm. Further insight may be obtained by using the dissipation scaling $\epsilon\propto k^{3/2}/L$, where $k$ is the turbulence kinetic energy, which is known to characterize homogeneous decaying turbulence far from initial conditions (Steiros, 2022a, b; Vassilicos, 2015). Expression (13) then becomes
@@
Inspection of the above expression reveals that our algorithm has, seemingly, collapsed the totality of cascade dynamics based on a single empirical length scale, $l\approx L^{0.38}\eta^{0.62}$, which is a combination of $L$ and $\eta$. There have been several attempts to develop single length scale theories of turbulence, derived by assuming self-similarity of the turbulence statistics, i.e., see von Karman & Howarth (1938), Sedov (2018), and the later theories of Barenblatt & Gavrilov (1974) and George (1992). The last two studies predict the Taylor microscale as the appropriate similarity variable, in close agreement with our data-driven methodology. In homogeneous decaying turbulence the Taylor microscale becomes $\lambda=L^{1/3}\eta^{2/3}$ (Pope, 2000), which is close to our empirically-derived length scale.
It might be tempting to interpret our empirical results as being in support of a single-scale picture for homogeneous turbulence, contrary to the commonly accepted view of turbulence as an intrinsically multi-length scale phenomenon (Batchelor, 1953; Townsend, 1976; Tennekes & Lumley, 1972; Pope, 2000). However, the results shown in figure 5 (and in most studies of homogeneous decaying turbulence) consider the far region of a turbulence grid where the integral length Reynolds number evolution is relatively slow, leading to a similarly slow evolution of $L/\eta$ (Vassilicos, 2015). As a result, our data shows that all length scales in question (integral, Kolmogorov, Taylor) produce a (visually) ‘good-enough’ collapse, as shown in figures 5(c)-(e), and inspection of any of these graphs might give the impression of a single-length scale self-similar evolution.
To better appreciate the single- or multi-length scale nature of the turbulence cascade, we need data for which the integral length scale Reynolds number varies drastically. To achieve this, we analyse the periodic box DNS data of decaying turbulence taken from Goto & Vassilicos (2016) (the details of the simulations can be found in the reference). Similar to the grid case, our goal is to collapse energy spectra taken at different decay times $\hat{t}$ (see figure 6(a)). Here, however, we consider datasets from two simulations which are identical in every aspect, apart from their size ($N=1024^{3}$ and $N=2048^{3}$) and the kinematic viscosity of the fluid, i.e., the Reynolds number of the flow varies drastically between the two cases, making the collapse of the curves less arbitrary.
Figure 6(a) presents the dimensional energy spectra from the two simulations, whereas figures 6(b) and 6(c) show the same data with the axes rescaled according to inertial and Kolmogorov similarity variables, respectively. The collapse of the large (small) scales is adequate only when the axes are normalized using the integral (Kolmogorov) scales, in agreement with a multiscale view of the cascade. However, data from the same simulation size retain a ‘good-enough’ collapse independent of the choice of normalization length, similar to our grid measurements, a fact which, if viewed in isolation, might give a false impression of a single length scale process. Figure 6(d) shows the algorithmically identified collapse of the data, that is, using the empirical transformations
$\tilde{\kappa}\propto\kappa L^{0.342}\eta^{0.658}$ (15a)
$\tilde{E_{11}}\propto E_{11}\epsilon^{-0.655}L^{-0.561}\eta^{-1.094}k^{-0.017}\approx E_{11}\epsilon^{-0.667}\left(L^{0.344}\eta^{0.656}\right)^{-5/3}.$ (15b)
The algorithm again identifies, approximately, the Taylor microscale ($\lambda=L^{1/3}\eta^{2/3})$ as the appropriate length for the collapse of the turbulence dynamics, but this time it cannot be claimed that the Taylor scale collapses both large and small-scale dynamics - only that it is an (algorithmically) optimal compromise when a single length scale collapse is enforced. We conjecture that the reason behind this could be explained from Lundgren’s two-scale turbulence theory which is based on the technique of matched asymptotic expansions (Lundgren, 2002, 2003). The repercussions of this theory are discussed in Obligado & Vassilicos (2019); Meldi & Vassilicos (2021): in particular, it is suggested that the Taylor microscale is the length scale at which the cascade dynamics are optimally distanced from both non-equilibrium effects (large scales) and dissipative effects (small scales) at the same time (in a matched asymptotic manner). The algorithmically identified collapse (figure 6(d)) produces best results in the intermediate, inertial range of the cascade where both above effects tend to become negligible. Based on this argument, the algorithm identifies the Taylor microscale scale because it is the optimal compromise of the two governing scales (inertial, Kolmogorov) of the cascade.
@@
Figure 6: Data-driven identification of self-similarity in decaying turbulence, using the DNS data of Goto & Vassilicos (2016). (a) Power spectral densities. (b) Spectrum normalized by the inertial scales. (c) Spectrum normalized by the Kolmogorov scales. (d) Spectrum normalized by the algorithmically identified expression (Eq. (15b)).
4 Discussion and conclusion
We have presented a methodology for identifying similarity variables from data, in the absence of governing equations. Given measurements of a quantity of interest, the method computes the discrete values of the similarity transformations that best collapse the data and expresses their analytic form via symbolic regression. The self-similarity that we identify is the optimal fit of the data, and is not derived from the governing partial differential equations and the boundary conditions of the specific problem.
The proposed method differs from dimensionless learning approaches that aim to reduce the dimensionality of parameter spaces via identification of dimensionless groups (Xie et al., 2022; Bakarji et al., 2022; Yuan & Lozano-Durán, 2025) because our focus is on discovering the similarity transformations that enable the collapse of entire solution profiles. Our approach is not restricted to dilational similarity transformations, and is capable of identifying more general, non-uniform relationships between variables, as demonstrated in the Burgers’ equation example. This flexibility broadens the range of problems for which similarity solutions can be uncovered. In practical terms, this approach also reduces the degree of prior knowledge required to reveal self-similar behaviour. For example, in the turbulent wake case, methods based on dimensional analysis would require the characteristic velocity difference to be provided as input in order to reveal the self-similarity. On the other hand, our method can uncover it autonomously. More generally, Step 1 of the proposed framework can be used to detect self-similar behaviour whilst being effectively agnostic to the underlying physics, i.e., without requiring prior specification of the relevant variables or parameters. In the symbolic regression step (Step 2), the user must supply the candidate parameters to obtain explicit analytic expressions, but the ability to first reveal the existence of self-similarity without specifying inputs may be particularly appealing for exploratory studies or complex systems. On the other hand, for cases of pure dilational self-similarity, methods based on dimensional analysis are expected to be more efficient. A note of caution: the computation of the $l_{2}$ norms in Eqn. (2) is performed over the common support of the transformed profiles. If the supports differ significantly, the optimization may become challenging. Although this issue did not arise in this work, we acknowledge it as a point that might need to be addressed in future work.
The capabilities of the method were demonstrated in five fluid mechanical problems, which are known or assumed to accept similarity solutions, based on data from both numerical and laboratory experiments. In four problems, the transformations identified by the method were in agreement to the analytically or empirically derived similarity transformations, whilst circumventing the need for scaling analysis or similarity arguments. In the fifth problem (turbulence cascade) our method identified an empirical length scale as the optimal similarity length of the cascade, which was close to the Taylor microscale. Further investigation revealed that this result cannot be treated as being unequivocally in support of single length scale theories of turbulence (von Karman & Howarth, 1938; Sedov, 2018; Barenblatt & Gavrilov, 1974; George, 1992), but rather expresses the role of the Taylor microscale as an intermediate scale between the integral and Kolmogorov scales, at which non-equilibrium and dissipative effects are optimally spaced (Meldi & Vassilicos, 2021).
The results illustrate ways that the proposed algorithm can be used for the identification of similarity and scaling laws in situations where rigorous mathematical analysis is challenging. This includes a wide range of applications in fluid physics but also in other processes such as quasicrystal shape and growth (Kamiya et al., 2018), stellar collapse (Yahil, 1983), single protein dynamics (Hu et al., 2016), and others. Other directions for future work include using the method to detect symmetry breaking, distinguish similarity across multiple scales, and identify the onset or breakdown of self-similarity.
[…]